OpenStudy (anonymous):

What is the square root of 676?

OpenStudy (anonymous):

26

sam (.sam.):

\[\huge \sqrt{676}=(676)^{\frac{1}{2}}=(26^{2})^{\frac{1}{2}}=26\]